### v3

# Evaluating Scoring Bias in LLM-as-a-Judge

Abstract
The remarkable capabilities of Large Language Models (LLMs) have led to the emergence of “LLM-as-a-Judge”, a paradigm where LLMs serve as automated evaluators for complex tasks. This approach has been widely adopted across diverse fields, including natural language processing and preference learning. However, the fairness and reliability of these LLM judges are often compromised by inherent biases. While current research has predominantly focused on biases in comparison-based evaluations, systematic investigations into scoring-based evaluations remain limited. To address this gap, we first formally define scoring bias as the discrepancy in scores when the judge model’s prompt is exposed to bias-inducing perturbations. We then propose a comprehensive framework to evaluate this bias. This framework is supported by two key contributions: (1) new evaluation datasets constructed by augmenting existing benchmarks with synthesized data, and (2) a suite of multi-faceted metrics designed for scoring bias analysis. Our experiments reveal that the scoring stability of LLM judges is significantly disrupted by these biases. Furthermore, our findings offer actionable insights for designing more robust scoring prompts and mitigating biases related to score rubrics, score IDs, and reference answer selection.
Introduction
Large Language Models (LLMs), with outstanding performance, have been widely applied to Natural Language Generation (NLG) tasks such as summarization (Zhang et al. 2024), question answering (Robinson, Rytting, and Wingate 2022), and domain-specific dialogues (Wang et al. 2024a; Singhal et al. 2025). In these tasks, the model’s responses require fine-grained evaluation of various aspects such as quality, factuality, and instruction-following, thereby providing valuable insights for model improvement. Traditional statistical NLG evaluation metrics, such as BLEU (Papineni et al. 2002) and ROUGE (Lin 2004), evaluate the generated text by measuring n-gram overlap between the text and the golden answer. The above methods exhibit two primary limitations: (1) They focus solely on the similarity between the generated text and the golden answer, rather than quantifying intrinsic attributes of the text such as writing style, factuality, and educational value. (2) The resulting metrics lack interpretability, making it difficult to provide guidance for the model improvement. There is an urgent need for fine-grained, multi-faceted, and interpretable evaluation metrics and methods tailored to NLG tasks and LLM applications. With remarkable capabilities in instruction understanding and response generation, LLMs can provide diverse forms of evaluation output, such as scoring, comparison, and ranking. In addition, they can conduct evaluations along specified dimensions and generate critiques for their judgments, thus enhancing the interpretability of the evaluation. This line of research, referred to as “LLM-as-a-Judge” (Gu et al. 2024; Li et al. 2024a), has recently emerged as a prominent focus.
Figure 1: Even slight perturbations to LLM judges lead to inconsistent scores for the same response.
Figure 2: Our evaluation framework to measure the scoring bias in LLM-as-a-Judge.
Robustness is a critical factor in ensuring the fairness and reliability of LLM judgments, which requires that, under identical judge rubrics, LLM judges should provide consistent judgments for semantically equivalent but differently phrased evaluation targets. However, recent work (Park et al. 2024) indicates that LLM judges are subject to various biases, which undermine the self-consistency of their judgments. Many studies have sought to identify the types of bias that cause shifts in judgments. Wang et al. finds that the judgments of the candidate responses can be easily influenced by altering their position in the context. Wataoka, Takahashi, and Ri measures the self-preference bias in LLM-as-a-Judge, which refers to the tendency of LLM judges to overestimate the quality of their own outputs. Wei et al. defines length bias as the tendency of LLM judges to favor longer or shorter responses.
Although existing studies have systematically investigated various biases in LLM-as-a-Judge, several pressing issues remain to be addressed. Firstly, while some work (Wang et al. 2024b; Ye et al. 2024a) has addressed biases such as positional bias and refinement awareness in scoring-based LLM judges, many existing studies remain focused on biases in comparative LLM judges. Existing work (Ye et al. 2024a) attempts to investigate specific biases in scoring tasks, such as self-enhancement bias and refinement-aware bias. However, there is still a lack of comprehensive research on biases in scoring-based LLM-as-a-Judge. Secondly, the biases defined in existing work are predominantly related to the choice of response models and the attributes or presentation forms of the evaluation targets, such as position and length. However, variation inherent in the LLM judge itself, such as the ordering of score rubrics and the selection of reference answers, may also impact the self-consistency of judgments, which lacks a systematic investigation in current research.
To solve the above problems, we propose scoring bias, defined as the scoring shifts of LLM-as-a-Judge under perturbations of the scoring prompt, as Figure 1 shows. In this work, we propose three types of scoring bias: score rubric order bias, score ID bias, and reference answer score bias, and design a framework to systematically evaluate these biases. We select four existing LLM-as-a-Judge benchmarks as evaluation datasets and utilize correlation with golden scores and score distributions as metrics to quantify the robustness and scoring tendency of LLM judges. Our key contributions can be summarized as follows:
  • We conduct the first systematic evaluation of scoring bias in LLM-as-a-Judge, introducing a formal definition of this bias and a comprehensive methodology to quantify it from multiple perspectives.
  • To address the critical lack of datasets for scoring bias analysis, we introduce an automatic data synthesis pipeline that expands existing LLM-as-a-Judge benchmarks into a tailored evaluation corpus, thereby establishing a foundational resource for future studies.
  • We empirically demonstrate that even the most advanced LLMs suffer from substantial scoring biases and exhibit obvious self-bias tendency. Furthermore, our analysis yields valuable recommendations for designing more robust scoring prompts and mitigating scoring biases.
Related Work
LLM-as-a-Judge
Gu et al. defines LLM-as-a-Judge as the task in which an LLM generates an evaluation result in any form given input data and context. Common forms of LLM judgment include score generation (Gao et al. 2023), pairwise comparison (Qin et al. 2024), ranking (Li, Patel, and Du 2023), multiple-choice selection (Li et al. 2024b), and critique generation (Ke et al. 2024). Improvement strategies for LLM-as-a-Judge include meticulous design of evaluation prompts (Liu et al. 2023), improvement of LLM’s abilities such as fine-tuning (Wang et al. 2024c), multi-LLM collaboration pipelines such as feedback mechanisms (Xu et al. 2023), and output optimization strategies such as multi-output ensembling (Sottana et al. 2023). The performance of LLM-as-a-Judge is usually measured using metrics that represent its alignment with human annotators or advanced LLM judgments such as agreement (Thakur et al. 2024), Cohen’s Kappa, and Spearman’s correlation (Bai et al. 2023). LLM-as-a-Judge is widely applied to evaluation (Lin and Chen 2023), alignment (Bai et al. 2022), retrieval (Sun et al. 2023), reasoning (Gao et al. 2024a), and other specific domains (Ryu et al. 2023; Babaei and Giudici 2024). LLM-as-a-Judge faces challenges of bias and vulnerability, which significantly compromise the fairness and reliability of LLM judges (Li et al. 2024a).
Bias in LLM-as-a-Judge
Ye et al. defines the prompt for LLM judges as three components: system instruction, question, and response. Based on this, they define bias in LLM-as-a-Judge as the tendency for judgments to differ when the system instruction or responses are modified in a bias-related manner. Common types of bias in LLM-as-a-Judge include position bias, length bias, self-enhancement bias, etc. These biases are often incorporated into the evaluation scope of existing LLM-as-a-Judge benchmarks. (Zheng et al. 2023; Wei et al. 2024) Currently, the types of bias remain a focal point of investigation, and numerous benchmarks have been proposed to evaluate bias in LLM-as-a-Judge. OffsetBias (Park et al. 2024) identifies six types of bias in LLM-as-a-Judge including length bias, concreteness bias, empty reference bias, and so on. CALM (Ye et al. 2024a) quantifies and analyzes 12 types of bias in LLM judges. However, the biases defined in current work are predominantly focused on comparative evaluation tasks, and there is a lack of systematic research and definitions specifically addressing biases in scoring-based evaluation tasks. Similar to research on adversarial attacks, the evaluation metrics for bias in LLM-as-a-Judge are often based on the consistency between bias-related modified judgments and the original judgments, or on alignment with golden answers.
Some studies focus on mitigating bias in LLM-as-a-Judge and aim to develop more robust LLM judges. Liu et al. introduces Pairwise-Preference Search (PAIRS) to mitigate biases in LLM evaluators. Wang et al. proposes three methods to calibrate positional bias in LLM-as-a-Judge, including multiple evidence calibration, balanced position calibration, and human-in-the-loop calibration. Our evaluation results on scoring bias will serve as a reference for future research on mitigating scoring biases in judge models and enhancing the robustness of scoring-based evaluations.
Evaluation Method
Figure 3: Scoring prompt template adapted from (Kim et al. 2024c), with red text indicating components that were perturbed in our experiments.
Figure 2 shows our evaluation framework, inspired by (Ye et al. 2024a). On a selected dataset, we inject the scoring prompt for the LLM judges with predefined perturbation types. Then we obtain the scores from a selected judge model. We evaluate the scoring bias by calculating the correlations between judge model’s scores and golden scores, which are obtained from advanced LLMs and human annotators. Furthermore, our metrics include score distribution, which is used to measure the impact of different types of bias on the scoring tendencies of LLM judges.
Dataset Data Size Human Judgment Reference Answer
BiGGen Bench 2,780
FLASK 2,001
MT Bench 320
Vicuna Bench 320
Table 1: Statistics of our evaluation datasets.
Problem Formulation
We assess the scoring bias in LLM-as-a-Judge by perturbing the scoring prompt and measuring the resulting impact on judgments. Following (Kim et al. 2024c), the input prompt $P$ for the scoring-based LLM judge consists of four components: a task description $T$, a target instance $C^{*}$, which can be either an instruction-response or a question-answer pair, a reference answer with a fixed score $R$, and score rubrics $S$, as shown in Figure 3. The input prompt $P$ is formulated as follows:
$P=(T,C^{*},[R],S),$ (1)
where * denotes the component will not be perturbed, and $\cdot$] means that the enclosed component is optional within the input prompt. This work focuses on biases induced by perturbations on the non-target components of the input prompt for LLM judges. Accordingly, we keep the target instance fixed. We also treat the reference answer as an optional element to explore its impact on scoring robustness, as it is not a mandatory part of the scoring process.
By perturbing the reference answer $R\rightarrow R^{\prime}$ or the score rubrics $S\rightarrow S^{\prime}$, we get a perturbed input prompt $P^{\prime}$. At the same time, the task description $T$ will also be modified to $T^{\prime}$ in conjunction with the perturbed components to ensure the prompt remains coherent. The score from the perturbed prompt will be:
$s^{\prime}=\textbf{LLM}(P^{\prime}).$ (2)
For each target instance $C$, we have a golden score $s$ from advanced LLMs or human annotators. If $s$ and $s^{\prime}$ differ, it demonstrates that the LLM judge exhibits scoring bias on the given sample.
Perturbation Types
We propose three methods to perturb key parts of the judge prompt to introduce scoring biases:
  • Score Rubric Order: The ordering of the description for each score within the score rubrics. e.g., ascending order (score 1 to score 5), descending order (score 5 to score 1), or random order.
  • Score ID: Typically, scores are represented using Arabic numerals $\{1,2,3,4,5\}$ as IDs. In this work, we explore the use of letter grades $\{\text{E},\text{D},\text{C},\text{B},\text{A}\}$ and Roman numerals $\{\text{i},\text{ii},\text{iii},\text{iv},\text{v}\}$ as scoring IDs.
  • Reference Answer Score: Based on the finding from (Kim et al. 2024a, c) that reference answers improve scoring accuracy for LLM judges, our work addresses an unexplored question: whether attaching specific scores to these references could degrade scoring robustness.
Datasets
Following previous work (Kim et al. 2024c; Chiang, Lee, and Lukasik 2025), we select four LLM-as-a-Judge datasets for scoring bias evaluation:
  • BiGGen Bench (Kim et al. 2024b): This is a principled generation benchmark with instance-specific evaluation criteria. We use the non-multilingual portion of its test set, which contains 2,780 responses for 695 instances, all evaluated by human annotators.
  • FLASK (Ye et al. 2024b): This is a fine-grained evaluation benchmark with 200 test prompts, 12 score rubrics, and 2,001 responses. It also includes scores labeled by human annotators.
  • MT Bench (Zheng et al. 2023): This is a multi-turn chat benchmark with 80 test prompts. Kim et al. write 80 hand-crafted score rubrics and generate 320 responses from WizardLM-13B (Xu et al. 2024), Vicuna-13B (Chiang et al. 2023), Llama-2-Chat-13B (Touvron et al. 2023), and GPT-3.5-Turbo-0613 (OpenAI 2023a).
  • Vicuna Bench (Chiang et al. 2023): This is a single-turn chat benchmark with 80 test prompts.Kim et al. also write 80 hand-crafted score rubrics and generate 320 responses from WizardLM-13B, Vicuna-13B, Llama-2-Chat-13B, and GPT-3.5-Turbo-0613.
Model BiGGen Bench FLASK
$\rho$ $r$ $\rho$ $r$
GPT-4.1 0.6048 0.6360 0.6401 0.6823
GPT-4o 0.6047 0.6340 0.6340 0.6557
GPT-4 0.5741 0.6045 0.6207 0.6862
Table 2: Comparison of the alignment between our LLM judgment pipeline and human golden scores, measured by Spearman’s $\rho$ and Pearson’s $r$ correlation coefficients. Scores for GPT-4.1 and GPT-4o are generated using our pipeline, whereas GPT-4 scores are sourced from the original benchmark.
The detailed statistics of our selected datasets are shown in Table 1. Due to the cost of manual annotation, the scores from advanced LLMs are usually used as golden scores (Ye et al. 2024b; Kim et al. 2024b). The instances in the above four datasets are pre-annotated with scores from GPT-4 (OpenAI 2023b). Due to the earlier release of GPT-4, we use more recent models GPT-4.1 (OpenAI 2025) and GPT-4o (OpenAI 2024) to obtain golden scores. Inspired by (Kim et al. 2024a, c), we repeat the scoring process three times using a judge prompt that includes a reference answer assigned a score of 5, and then take a majority vote to obtain the final score from the LLM evaluators. Table 2 shows that the GPT-4.1 scores exhibit the highest correlation with human scores. Throughout the remainder of our experiments, the scores generated by GPT-4.1 are treated as the golden LLM scores.
Each instance in BiGGen Bench and FLASK is accompanied by a reference answer with a score of 5. To measure the impact of the reference answer score on the judgment robustness, it is necessary to additionally generate reference answers with scores ranging from 1 to 4. We propose a generation-review pipeline to synthesize reference answers for each sample in FLASK, as shown in Figure 4. Given an instruction and score rubrics, the response model generates the response according to the description from score 1 to 4. Then a review model checks whether the response matches the corresponding score description. If they match, we accept the generated responses; otherwise, we rerun the pipeline until a match is achieved. In this pipeline, we use GPT-4.1 and GPT-4o, alternating their roles as the response and review models.
Metrics
Figure 4: The proposed generation-review pipeline for synthesizing reference answers.
We quantify the effects of scoring bias on both scoring accuracy and scoring tendency. Given the golden scores, $S=\{s_{1},\cdots,s_{n}\}$, and the LLM judge’s scores, $S^{\prime}=\{s^{\prime}_{1},\cdots,s^{\prime}_{n}\}$, we quantify scoring accuracy as the correlation between $S$ and $S^{\prime}$. Specifically, we report both Spearman’s correlation coefficient $\rho$ (Spearman 1904) and Pearson’s correlation coefficient $r$ (Pearson 1895). The scoring tendency is quantified by the score distribution:
$N_{k}$ $=\sum_{i=1}^{n}{\mathbb{I}(s_{i}=k)},\quad k\in\{1,2,3,4,5\},$ (3)
$N$ $=(N_{1},N_{2},N_{3},N_{4},N_{5}).$ (4)
Here, $N_{k}$ represents the count of score $k$ in the scoring results, and $N$ denotes the distribution of the score counts.
Experiments
Model Perturbation Strategy BiGGen Bench FLASK MT Bench Vicuna Bench
Human GPT-4.1 Human GPT-4.1 GPT-4.1 GPT-4.1
$\rho$ $r$ $\rho$ $r$ $\rho$ $r$ $\rho$ $r$ $\rho$ $r$ $\rho$ $r$
GPT-4o Baseline - 0.5685 0.6165 0.7572 0.8059 0.5529 0.5914 0.7369 0.7676 0.7578 0.7724 0.6862 0.7696
Rubrics Order Descending 0.5852 0.6245 0.7787 0.8233 0.5754 0.6080 0.7602 0.7760 0.7479 0.7597 0.6734 0.7635
Random 0.5963 0.6313 0.7698 0.8124 0.5592 0.5893 0.7420 0.7573 0.7673 0.7838 0.6701 0.7504
Score IDs Letter 0.5641 0.6034 0.7595 0.7994 0.5558 0.5977 0.7341 0.7609 0.7317 0.7454 0.6539 0.7712
Roman numerals 0.5754 0.6171 0.7709 0.8148 0.5535 0.5912 0.7380 0.7641 0.7725 0.7786 0.7073 0.7821
DeepSeek-V3 Baseline - 0.5314 0.5748 0.6775 0.7394 0.5163 0.5660 0.6711 0.7063 0.6323 0.6593 0.4313 0.5443
Rubrics Order Descending 0.5356 0.5812 0.6611 0.7211 0.5181 0.5554 0.6555 0.6858 0.6183 0.6436 0.4993 0.5854
Random 0.5136 0.5609 0.6547 0.7125 0.4999 0.5470 0.6521 0.6752 0.6178 0.6460 0.5086 0.5482
Score IDs Letter 0.5378 0.5689 0.6803 0.7179 0.5314 0.5755 0.6663 0.6999 0.6489 0.6715 0.4832 0.5657
Roman numerals 0.5164 0.5636 0.6688 0.7195 0.5066 0.5542 0.6750 0.7109 0.6558 0.6642 0.4416 0.5285
Qwen3-32B Baseline - 0.5421 0.5855 0.7328 0.7715 0.5323 0.5695 0.6549 0.6807 0.6959 0.7053 0.6419 0.7188
Rubrics Order Descending 0.5572 0.5989 0.7443 0.7862 0.5345 0.5657 0.6586 0.6839 0.6843 0.6989 0.6205 0.7020
Random 0.5528 0.5938 0.7404 0.7824 0.5115 0.5456 0.6537 0.6748 0.6554 0.6688 0.6416 0.7252
Score IDs Letter 0.5353 0.5851 0.7280 0.7770 0.5393 0.5724 0.6650 0.6975 0.6795 0.6930 0.6157 0.6808
Roman numerals 0.5139 0.5629 0.6992 0.7412 0.5257 0.5781 0.6630 0.6980 0.6360 0.6418 0.6072 0.7126
Qwen3-8B Baseline - 0.5092 0.5308 0.6522 0.6492 0.5035 0.5309 0.6464 0.6616 0.6238 0.6342 0.6639 0.7284
Rubrics Order Descending 0.5241 0.5456 0.6458 0.6559 0.4885 0.5339 0.6315 0.6566 0.6005 0.6135 0.5500 0.6647
Random 0.4869 0.5025 0.6091 0.5950 0.4778 0.5004 0.6061 0.6216 0.5949 0.5983 0.5135 0.6053
Score IDs Letter 0.4853 0.5117 0.6410 0.6665 0.4873 0.5244 0.6334 0.6586 0.5985 0.6126 0.5812 0.6890
Roman numerals 0.4801 0.5120 0.6231 0.6445 0.4606 0.5087 0.6177 0.6468 0.6016 0.5975 0.6111 0.7178
Mistral-24B Baseline - 0.5059 0.5495 0.6910 0.7201 0.4950 0.5385 0.6303 0.6593 0.5943 0.5762 0.6116 0.6917
Rubrics Order Descending 0.5369 0.5761 0.6987 0.7213 0.4922 0.5327 0.6362 0.6534 0.6104 0.6102 0.6137 0.7082
Random 0.5053 0.5450 0.6569 0.6841 0.4756 0.5042 0.6212 0.6338 0.5603 0.5602 0.5461 0.6418
Score IDs Letter 0.5114 0.5632 0.6685 0.7084 0.4880 0.5339 0.6311 0.6594 0.5375 0.5485 0.5558 0.6608
Roman numerals 0.5102 0.5590 0.6640 0.6962 0.4957 0.5494 0.6533 0.6872 0.5863 0.5907 0.5903 0.6905
Table 3: Correlation between golden scores and judge model scores under various prompt perturbations. For each judge model, the first row (Baseline) presents the metrics for the original, unperturbed prompt. The subsequent rows show the results under different types of perturbations applied to the rubrics order and score IDs. The degradation in correlation scores compared to the baseline reflects the model’s scoring bias, indicating its lack of robustness to these perturbations. Model names are abbreviated: DeepSeek-V3 denotes DeepSeek-V3-67B and Mistral-24B denotes Mistral-Small-24B-Instruct. We report Spearman’s ($\rho$) and Pearson’s ($r$) correlation coefficient.
Setup
Judge Models. Previous work (Gao et al. 2024b; Li et al. 2024c; Liu et al. 2024b) has reached a consensus that judge models should be stronger models to ensure evaluation reliability (Ye et al. 2024a). After preliminary experiments, we found that scoring bias is a common phenomenon in judge models based on various LLMs. Furthermore, we aim to explore the impact of factors such as model type and size on scoring bias. Therefore, we select diverse judge models that include closed-source, open-source, large-size, and small-size models. They are GPT-4o, DeepSeek-V3-671B (Liu et al. 2024a), Qwen3-32B, Qwen3-8B (Yang et al. 2025), and Mistral-Small-24B-Instruct-2501 (MistralAI 2025). These judge models we selected differ from the models used to generate responses in the evaluation datasets, thus mitigating potential self-enhancement bias (Zheng et al. 2023; Ye et al. 2024a).
Judge Prompt. The prompt we use for scoring-based judgment, adapted from (Kim et al. 2024c), is shown in Figure 3. The unperturbed prompt is the one with ascending score rubrics, Arabic numeral score IDs, and no reference answer.
Hyper-parameters. To ensure a deterministic scoring process, which is crucial for a fair evaluation of bias, we set the temperature of the LLM judge to 0 during all scoring tasks.
Main Results
Figure 5: Spearman’s correlation between human annotator scores and judge model scores on the FLASK dataset, under perturbations applied to the reference answer scores.
The existence of scoring bias is proved in all judge models. Table 3 and Figure 5 show the correlation between judge model scores and golden scores under all perturbation methods. For all judge models we evaluated, including advanced models such as GPT-4o, we observed fluctuations in the scoring results under different types of perturbations. This indicates that scoring bias significantly impacts the robustness of LLM-as-a-Judge. Furthermore, different judge models exhibit varying sensitivity to scoring bias:
(1) More powerful judge models tend to be more robust to scoring bias. Consistent with existing consensus (Gao et al. 2024b; Li et al. 2024c; Liu et al. 2024b), our experimental results indicate that judge models with advanced capability and larger size tend to exhibit higher reliability in scoring. As shown in Table 3, for example, across all perturbation methods and datasets, the fluctuation in correlation coefficients between GPT-4o and the golden scores remains within 0.03, except for the case involving a reference answer with a score of 5. A comparison of the results for Qwen3-32B and Qwen3-8B shows that judge models with larger parameter sizes demonstrate greater robustness in scoring. In comparison with the GPT-4.1 scores in BiGGen-Bench and Vicuna Bench, the correlation coefficients of the Qwen3-8B scores exhibit maximum fluctuations exceeding 0.05 under score rubric order bias and score ID bias. With reference answer score bias, the maximum fluctuation reaches nearly 0.2. However, the changes in the correlation coefficient of Qwen3-32B under biases are generally within 0.03. Therefore, more capable judge models should be prioritized in scoring evaluations.
(2) Contrary to the conventional view that bias is always harmful, we find that scoring bias in LLM-as-a-Judge can also have surprisingly positive effects on the evaluation results. For example, as Table 3 shows, in most cases, the correlation between GPT-4o scores and golden scores increases when Roman numerals are used as scoring IDs. However, when DeepSeek-V3-671B, Qwen3-32B and Qwen3-8B are employed as judge models, the use of Roman numerals generally exerts a negative effect on their scoring accuracy. An interesting phenomenon is that the impact of reference answer score bias on the judge models we evaluated appears to be consistent, which will be further analyzed in subsequent sections. Our experimental results can serve as a valuable reference for the selection of judge models and the design of scoring prompt templates.
(3) Judge models have their own scoring tendencies. Figure 6 presents the scoring tendencies of different judge models under various scoring biases on the FLASK dataset in the form of heatmaps. It is obvious that each judge model exhibits its own specific scoring preferences. Under unbiased conditions, DeepSeek-V3-671B assigns a score of 5 to more than half of the instruction-response pairs. Compared to the human scoring distribution, Mistral-Small-24B-Instruct has a stronger tendency to assign a score of 4 to the samples. Other judge models show smoother scoring distribution shifts, but certain scores are still preferred. GPT-4o demonstrates a preference for a score of 4, while Qwen3-32B favors a score of 5. An examination of Figure 6 reveals that score rubric order bias and score ID bias have a certain impact on the scoring tendencies of the judge models; however, they do not fundamentally alter the overall score distribution. An exception is Qwen3-8B, whose score distribution is more significantly influenced by biases, possibly due to its small size. It is particularly evident that the reference answer score bias exerts a substantial impact on the score distribution, which will be analyzed in detail in the following sections. Our experimental results show that judge models inherently exhibit biases in their scoring tendencies. Therefore, debiasing the judge models is crucial to ensuring the fairness and reliability of the evaluation.
Figure 6: Score distributions of LLM judges under various prompt perturbations on the FLASK dataset. The horizontal axis represents the assigned score (where -1 indicates a failed judgment), and the vertical axis lists the different perturbation methods applied. For reference, the top row in each heatmap (Human) shows the distribution of human scores. The second row (None) shows the model’s score distribution for the original, unperturbed prompt.
Methods for Scoring Bias Mitigation
For advanced or large size judge models, applying scoring prompt that deviate from conventional preferences is worth considering.
In almost all scoring-based evaluation tasks, Arabic numerals are commonly used as scoring IDs. The score rubric order typically follows a sequential order from score 1 to score 5. These choices align with human writing conventions and cognitive preferences in the design of scoring prompts. However, our experimental results shown in Table 3 indicate that the scoring prompts designed to align with the usual preferences do not necessarily produce the most accurate scoring results. As mentioned earlier, the performance of the GPT-4o judge model in scoring evaluations improves when Roman numerals are used as scoring IDs. In more than half of the cases, using letter grades as scoring IDs improves the scoring accuracy of DeepSeek-V3-671B. Regarding the score rubric order bias, adopting a descending order improves the scoring accuracy of GPT-4o, Qwen3-32B, and Mistral-Small-24B-Instruct in most instances. Intuitively, employing a random scoring rubric order tends to negatively impact the scoring accuracy of all judge models. Based on our analysis, we propose a core design principle for scoring prompts: intentional yet meaningful deviation from convention. Effective strategies include introducing structured, unconventional formats, such as presenting scoring rubrics in a descending order or using non-numeric identifiers like letter grades (A, B, C). However, the key is to maintain clarity and logical coherence. Therefore, chaotic perturbations—such as a fully randomized rubric order or non-hierarchical identifiers—should be avoided. Moreover, inspired by previous work (Liu et al. 2024c; Park et al. 2024), perturbations applied to scoring prompt also show potential as debiasing strategies. For example, effective approaches include implementing multiple variations of the scoring rubric order or IDs and then aggregating the results, or adversarially training the models on samples enhanced with scoring bias perturbations.
Using full-mark reference answers can stabilize scoring preferences and achieve optimal accuracy.
The impact of reference answer score bias on the scoring accuracy and scoring tendency of the judge model is presented in Figure 5 and Figure 6. The experimental results indicate that, compared to the score rubric order bias and the score ID bias, the influence of the reference answer score bias on scoring robustness is both more pronounced and more regular. Figure 5 illustrates the impact of introducing reference answers with different scores in the input prompt on the scoring accuracy of the judge model. For GPT-4o, DeepSeek-V3-671B, and Mistral-Small-24B-Instruct, the scoring accuracy in most cases improves when reference answers are introduced. However, for the two Qwen judge models, only reference answers with scores of 4 or 5 contribute to enhancing scoring performance. Figure 5 also reveals an interesting phenomenon: the higher the score of the reference answer, the better the scoring accuracy of the judge model. Although DeepSeek-V3-671B, Qwen3-32B, and Mistral-Small-24B-Instruct do not strictly adhere to this pattern, the correlation between their scores and human scores similarly exhibits an upward trend as the reference answer score increases.
Figure 6 shows that the score of the reference answer has a significant impact on the scoring tendency of the judge model. Specifically, the scoring results from DeepSeek-V3-671B, Qwen3-32B, and Qwen3-8B appear to be highly influenced by the score of the reference answer. When the score of the reference answer is between 1 and 4, these models tend to assign scores similar to that of the reference answer. Interestingly, when a reference answer with a score of 5 is introduced, the scoring behavior of the judge model appears to become more rational. This is particularly notable for DeepSeek-V3-671B, which exhibits a strong tendency to assign a score of 5 without bias. However, the introduction of a reference answer scored at 5 mitigates this phenomenon.
The above analysis reveals that the score of the reference answer introduced in the prompt has a significant impact on the judge model’s scoring behavior. In particular, when reference answers with a score of 5 are introduced, the models achieve the highest scoring accuracy without disproportionately influencing their scoring preferences. Therefore, it can be concluded that when incorporating reference answers into a scoring-based judge prompt, a full-mark reference answer is the best choice. Interestingly, this conceptually aligns with the principles of traditional statistical metrics, such as BLEU and ROUGE, which emphasize the critical role of golden answers in evaluation.
Limitations and Future Work
While our work offers notable conclusions and insights, several limitations remain to be addressed in future work.
Firstly, our evaluation framework has room for further expansion. More types of scoring biases remain to be identified. Building more comprehensive benchmarks to evaluate scoring bias in LLM-as-a-Judge is also warranted, as a limited data distribution may lead to an unfair evaluation.
Secondly, mitigating scoring bias needs further exploration. Although there are no dedicated methods for addressing scoring bias, intuitive and straightforward approaches, such as scoring multiple times followed by majority voting or averaging the scores, require empirical validation for their effectiveness in mitigating scoring bias. Moreover, the elaborate design of scoring prompt templates, the construction of debiased scoring-based judgment datasets (Park et al. 2024), and adversarial training are potential approaches to enhancing the scoring robustness of judge models.
Additionally, although our work demonstrates the existence of scoring bias in LLM-as-a-Judge, the underlying causes of scoring bias remain to be validated. Approaches such as training data analysis and information flow observation (Yuan et al. 2025) may help identify the reasons for scoring bias, whether it originates from within the model or from external factors.
Conclusion
This work investigates scoring bias in LLM-as-a-Judge, where scoring judgments shift due to minor perturbations in judge prompts. To systematically evaluate this, we designed a framework that perturbs score rubric order, score IDs, and reference answers, quantifying the bias using scoring accuracy and scoring tendency. Experiments on synthesized datasets expanded from four benchmarks confirm that scoring bias significantly undermines the robustness and reliability of existing LLM judges, with varying sensitivity across different models. To mitigate this, we recommend unconventional prompt designs—such as descending rubric order or using letter or Roman numeral IDs—and providing a full-mark reference answer when applicable, both of which effectively improve performance. In summary, our work provides systematic evidence and actionable recommendations for evaluating and enhancing the reliability of LLM judges.

### v4

@@
 Abstract
-The remarkable capabilities of Large Language Models (LLMs) have led to the emergence of “LLM-as-a-Judge”, a paradigm where LLMs serve as automated evaluators for complex tasks.
-This approach has been widely adopted across diverse fields, including natural language processing and preference learning.
-However, the fairness and reliability of these LLM judges are often compromised by inherent biases.
-While current research has predominantly focused on biases in comparison-based evaluations, systematic investigations into scoring-based evaluations remain limited.
-To address this gap, we first formally define scoring bias as the discrepancy in scores when the judge model’s prompt is exposed to bias-inducing perturbations.
-We then propose a comprehensive framework to evaluate this bias.
-This framework is supported by two key contributions: (1) new evaluation datasets constructed by augmenting existing benchmarks with synthesized data, and (2) a suite of multi-faceted metrics designed for scoring bias analysis.
-Our experiments reveal that the scoring stability of LLM judges is significantly disrupted by these biases.
-Furthermore, our findings offer actionable insights for designing more robust scoring prompts and mitigating biases related to score rubrics, score IDs, and reference answer selection.
-Introduction
-Large Language Models (LLMs), with outstanding performance, have been widely applied to Natural Language Generation (NLG) tasks such as summarization (Zhang et al.
-2024), question answering (Robinson, Rytting, and Wingate 2022), and domain-specific dialogues (Wang et al.
-2024a; Singhal et al.
-2025).
-In these tasks, the model’s responses require fine-grained evaluation of various aspects such as quality, factuality, and instruction-following, thereby providing valuable insights for model improvement.
-Traditional statistical NLG evaluation metrics, such as BLEU (Papineni et al.
-2002) and ROUGE (Lin 2004), evaluate the generated text by measuring n-gram overlap between the text and the golden answer.
-The above methods exhibit two primary limitations: (1) They focus solely on the similarity between the generated text and the golden answer, rather than quantifying intrinsic attributes of the text such as writing style, factuality, and educational value.
-(2) The resulting metrics lack interpretability, making it difficult to provide guidance for the model improvement.
-There is an urgent need for fine-grained, multi-faceted, and interpretable evaluation metrics and methods tailored to NLG tasks and LLM applications.
-With remarkable capabilities in instruction understanding and response generation, LLMs can provide diverse forms of evaluation output, such as scoring, comparison, and ranking.
-In addition, they can conduct evaluations along specified dimensions and generate critiques for their judgments, thus enhancing the interpretability of the evaluation.
-This line of research, referred to as “LLM-as-a-Judge” (Gu et al.
-2024; Li et al.
-2024a), has recently emerged as a prominent focus.
+The “LLM-as-a-Judge” paradigm, using Large Language Models (LLMs) as automated evaluators, is pivotal to LLM development, offering scalable feedback for complex tasks.
+However, the reliability of these judges is compromised by various biases.
+Existing research has heavily concentrated on biases in comparative evaluations.
+In contrast, scoring-based evaluations—which assign an absolute score and are often more practical in industrial applications—remain under-investigated.
+To address this gap, we undertake the first dedicated examination of scoring bias in LLM judges.
+We shift the focus from biases tied to the evaluation targets to those originating from the scoring prompt itself.
+We formally define scoring bias and identify three novel, previously unstudied types: rubric order bias, score ID bias, and reference answer score bias.
+We propose a comprehensive framework to quantify these biases, featuring a suite of multi-faceted metrics and an automatic data synthesis pipeline to create a tailored evaluation corpus1.
+Our experiments empirically demonstrate that even the most advanced LLMs suffer from these substantial scoring biases.
+Our analysis yields actionable insights for designing more robust scoring prompts and mitigating these newly identified biases.
+footnotetext: Equal contribution. Work done while at Ant Groupfootnotetext: 1Our dataset is publicly available at https://github.com/KMdsy/scoring_bias/
+1 Introduction
+The “LLM-as-a-Judge” [9, 15] paradigm, where Large Language Models (LLMs) serve as automated evaluators, is pivotal to LLM development.
+These judges provide fine-grained assessments for complex tasks like factuality evaluation [32] and instruction-following [50], and act as scalable proxies for human feedback in preference learning [2, 14].
+The increasing prominence of this paradigm makes its robustness a critical concern.
+Ideally, LLM judges should provide objective judgments, uninfluenced by irrelevant factors.
+However, recent work [29] confirms that LLM judges exhibit various biases, undermining their consistency.
+Existing research has identified several such biases, including position bias [40], self-preference bias [42], length bias [43], and so on.
 Figure 1: Even slight perturbations to LLM judges lead to inconsistent scores for the same response.
-Figure 2: Our evaluation framework to measure the scoring bias in LLM-as-a-Judge.
-Robustness is a critical factor in ensuring the fairness and reliability of LLM judgments, which requires that, under identical judge rubrics, LLM judges should provide consistent judgments for semantically equivalent but differently phrased evaluation targets.
-However, recent work (Park et al.
-2024) indicates that LLM judges are subject to various biases, which undermine the self-consistency of their judgments.
-Many studies have sought to identify the types of bias that cause shifts in judgments.
-Wang et al.
-finds that the judgments of the candidate responses can be easily influenced by altering their position in the context.
-Wataoka, Takahashi, and Ri measures the self-preference bias in LLM-as-a-Judge, which refers to the tendency of LLM judges to overestimate the quality of their own outputs.
-Wei et al.
-defines length bias as the tendency of LLM judges to favor longer or shorter responses.
-Although existing studies have systematically investigated various biases in LLM-as-a-Judge, several pressing issues remain to be addressed.
-Firstly, while some work (Wang et al.
-2024b; Ye et al.
-2024a) has addressed biases such as positional bias and refinement awareness in scoring-based LLM judges, many existing studies remain focused on biases in comparative LLM judges.
-Existing work (Ye et al.
-2024a) attempts to investigate specific biases in scoring tasks, such as self-enhancement bias and refinement-aware bias.
-However, there is still a lack of comprehensive research on biases in scoring-based LLM-as-a-Judge.
-Secondly, the biases defined in existing work are predominantly related to the choice of response models and the attributes or presentation forms of the evaluation targets, such as position and length.
-However, variation inherent in the LLM judge itself, such as the ordering of score rubrics and the selection of reference answers, may also impact the self-consistency of judgments, which lacks a systematic investigation in current research.
-To solve the above problems, we propose scoring bias, defined as the scoring shifts of LLM-as-a-Judge under perturbations of the scoring prompt, as Figure 1 shows.
-In this work, we propose three types of scoring bias: score rubric order bias, score ID bias, and reference answer score bias, and design a framework to systematically evaluate these biases.
-We select four existing LLM-as-a-Judge benchmarks as evaluation datasets and utilize correlation with golden scores and score distributions as metrics to quantify the robustness and scoring tendency of LLM judges.
+While existing studies have investigated these biases, pressing issues remain.
+Research has heavily concentrated on comparative LLM judges, which determine the superior of two responses.
+However, scoring-based evaluation, providing an absolute quality score for a single response, is often more practical and widely adopted.
+It is more intuitive and better suited for standardized, quantitative analysis, especially when a direct comparison object is unavailable [39].
+Consequently, an in-depth investigation into the biases of scoring-based judges is crucial.
+Research in this area remains unsystematic; while some issues like self-enhancement bias, refinement-aware bias [47], and judge model choice bias [40] have been identified, the exploration is far from comprehensive and leaves many potential biases unaddressed.
+This research gap is not merely theoretical.
+Our work is motivated by a practical observation: when using a score-based LLM judge to evaluate a financial advisory assistant, we found that adding a single in-context example for a specific score level (e.g., an exemplar for a 4-point response) caused the absolute scores of other, unrelated samples to shift.
+This instability highlights a critical, unaddressed sensitivity to prompt examples and served as the primary catalyst for our investigation into other potential biases affecting score-based judges.
+In this work, we consider the scoring bias as the measurable scoring shifts exhibited by an LLM-as-a-Judge when subjected to perturbations to its scoring prompt (as illustrated in Figure 1).
+Diverging from prior work, which has focused heavily on biases tied to the evaluation targets, our research investigates biases originating from other components of the scoring prompt.
+Specifically, we identify and define three novel and previously unstudied types of scoring bias: rubric order bias, score ID bias, and reference answer score bias.
+To identify and quantify these three biases, we propose a comprehensive evaluation framework.
+This framework is designed to be extensible, enabling the identification of various scoring biases beyond those explored in this paper.
+We conduct experiments on four existing LLM-as-a-Judge benchmarks and utilize a suite of metrics, categorized as stability metrics, accuracy metrics, and scoring tendency, to thoroughly quantify the impact of these biases.
 Our key contributions can be summarized as follows:
-  • We conduct the first systematic evaluation of scoring bias in LLM-as-a-Judge, introducing a formal definition of this bias and a comprehensive methodology to quantify it from multiple perspectives.
-  • To address the critical lack of datasets for scoring bias analysis, we introduce an automatic data synthesis pipeline that expands existing LLM-as-a-Judge benchmarks into a tailored evaluation corpus, thereby establishing a foundational resource for future studies.
-  • We empirically demonstrate that even the most advanced LLMs suffer from substantial scoring biases and exhibit obvious self-bias tendency.
-Furthermore, our analysis yields valuable recommendations for designing more robust scoring prompts and mitigating scoring biases.
-Related Work
-LLM-as-a-Judge
-Gu et al.
-defines LLM-as-a-Judge as the task in which an LLM generates an evaluation result in any form given input data and context.
-Common forms of LLM judgment include score generation (Gao et al.
-2023), pairwise comparison (Qin et al.
-2024), ranking (Li, Patel, and Du 2023), multiple-choice selection (Li et al.
-2024b), and critique generation (Ke et al.
-2024).
-Improvement strategies for LLM-as-a-Judge include meticulous design of evaluation prompts (Liu et al.
-2023), improvement of LLM’s abilities such as fine-tuning (Wang et al.
-2024c), multi-LLM collaboration pipelines such as feedback mechanisms (Xu et al.
-2023), and output optimization strategies such as multi-output ensembling (Sottana et al.
-2023).
-The performance of LLM-as-a-Judge is usually measured using metrics that represent its alignment with human annotators or advanced LLM judgments such as agreement (Thakur et al.
-2024), Cohen’s Kappa, and Spearman’s correlation (Bai et al.
-2023).
-LLM-as-a-Judge is widely applied to evaluation (Lin and Chen 2023), alignment (Bai et al.
-2022), retrieval (Sun et al.
-2023), reasoning (Gao et al.
-2024a), and other specific domains (Ryu et al.
-2023; Babaei and Giudici 2024).
-LLM-as-a-Judge faces challenges of bias and vulnerability, which significantly compromise the fairness and reliability of LLM judges (Li et al.
-2024a).
-Bias in LLM-as-a-Judge
-Ye et al.
-defines the prompt for LLM judges as three components: system instruction, question, and response.
+  • We identify three novel types of scoring bias: rubric order bias, score ID bias, and reference answer score bias.
+To facilitate their evaluation, we introduce an automatic data synthesis pipeline that expands existing benchmarks into a tailored corpus, establishing a foundational resource for future research.
+  • We propose the first evaluation framework for investigating and quantifying scoring bias, introducing metrics across three key dimensions: stability, accuracy, and scoring tendency.
+  • We empirically demonstrate that even the most advanced LLMs suffer from the substantial scoring biases we identify.
+Furthermore, our analysis yields valuable recommendations for designing more robust scoring prompts and mitigating these biases.
+2 Related Work
+2.1 LLM-as-a-Judge
+Previous work [9] defines LLM-as-a-Judge as the task in which an LLM generates an evaluation result in any form given input data and context.
+Common forms of LLM judgment include score generation [8], pairwise comparison [31], ranking [17], multiple-choice selection [16], and critique generation [10].
+Improvement strategies for LLM-as-a-Judge include meticulous design of evaluation prompts [22], improvement of LLM’s abilities such as fine-tuning [41], multi-LLM collaboration pipelines such as feedback mechanisms [45], and output optimization strategies such as multi-output ensembling [34].
+The performance of LLM-as-a-Judge is usually measured using metrics that represent its alignment with human annotators or advanced LLM judgments such as agreement [37], Cohen’s Kappa, and Spearman’s correlation [3].
+LLM-as-a-Judge is widely applied to evaluation [19], alignment [2], retrieval [36], reasoning [6], and other specific domains [1, 33].
+LLM-as-a-Judge faces challenges of bias and vulnerability, which significantly compromise the fairness and reliability of LLM judges [15].
+2.2 Bias in LLM-as-a-Judge
+Existing work [47] defines the prompt for LLM judges as three components: system instruction, question, and response.
 Based on this, they define bias in LLM-as-a-Judge as the tendency for judgments to differ when the system instruction or responses are modified in a bias-related manner.
 Common types of bias in LLM-as-a-Judge include position bias, length bias, self-enhancement bias, etc.
-These biases are often incorporated into the evaluation scope of existing LLM-as-a-Judge benchmarks.
-(Zheng et al.
-2023; Wei et al.
-2024) Currently, the types of bias remain a focal point of investigation, and numerous benchmarks have been proposed to evaluate bias in LLM-as-a-Judge.
-OffsetBias (Park et al.
-2024) identifies six types of bias in LLM-as-a-Judge including length bias, concreteness bias, empty reference bias, and so on.
-CALM (Ye et al.
-2024a) quantifies and analyzes 12 types of bias in LLM judges.
+These biases are often incorporated into the evaluation scope of existing LLM-as-a-Judge benchmarks [43, 51].
+Currently, the types of bias remain a focal point of investigation, and numerous benchmarks have been proposed to evaluate bias in LLM-as-a-Judge.
+OffsetBias [29] identifies six types of bias in LLM-as-a-Judge including length bias, concreteness bias, empty reference bias, and so on.
+CALM [47] quantifies and analyzes 12 types of bias in LLM judges.
 However, the biases defined in current work are predominantly focused on comparative evaluation tasks, and there is a lack of systematic research and definitions specifically addressing biases in scoring-based evaluation tasks.
@@
 Some studies focus on mitigating bias in LLM-as-a-Judge and aim to develop more robust LLM judges.
-Liu et al.
-introduces Pairwise-Preference Search (PAIRS) to mitigate biases in LLM evaluators.
-Wang et al.
-proposes three methods to calibrate positional bias in LLM-as-a-Judge, including multiple evidence calibration, balanced position calibration, and human-in-the-loop calibration.
+[23] introduces Pairwise-Preference Search (PAIRS) to mitigate biases in LLM evaluators.
+[40] proposes three methods to calibrate positional bias, including multiple evidence calibration, balanced position calibration, and human-in-the-loop calibration.
 Our evaluation results on scoring bias will serve as a reference for future research on mitigating scoring biases in judge models and enhancing the robustness of scoring-based evaluations.
-Evaluation Method
-Figure 3: Scoring prompt template adapted from (Kim et al. 2024c), with red text indicating components that were perturbed in our experiments.
-Figure 2 shows our evaluation framework, inspired by (Ye et al.
-2024a).
-On a selected dataset, we inject the scoring prompt for the LLM judges with predefined perturbation types.
+3 Evaluation Method
+Figure 2: Our evaluation framework to measure the scoring bias in LLM-as-a-Judge.
+Figure 2 shows our evaluation framework, inspired by [47].
+On a selected dataset, we inject predefined perturbation types into the scoring prompts for the LLM judges.
 Then we obtain the scores from a selected judge model.
-We evaluate the scoring bias by calculating the correlations between judge model’s scores and golden scores, which are obtained from advanced LLMs and human annotators.
+We evaluate the scoring bias by calculating the consistency between scores before and after injecting perturbations, and the correlations between the judge model’s scores and the golden scores (obtained from advanced LLMs and human annotators).
 Furthermore, our metrics include score distribution, which is used to measure the impact of different types of bias on the scoring tendencies of LLM judges.
+3.1 Problem Formulation
+Scoring Prompt ### Task Description: An instruction (might include an Input inside it), a response to evaluate, a reference answer that gets a score of 5, and a score rubric representing a evaluation criteria are given.
+1.
+Write a detailed feedback that assess the quality of the response strictly based on the given score rubric, not evaluating in general.
+2.
+After writing a feedback, write a score that is an integer between 1 and 5.
+You should refer to the score rubric.
+3.
+The output format should look as follows: “Feedback: (write a feedback for criteria) [RESULT] (an integer number between 1 and 5) 4.
+Please do not generate any other opening, closing, and explanations.
+### Reference Answer (Score 5): {orig_reference_answer} ### Score Rubrics: [{score_criteria}] Score 1: {score1_description} Score 2: {score2_description} Score 3: {score3_description} Score 4: {score4_description} Score 5: {score5_description} ### The instruction to evaluate: {orig_instruction} ### Response to evaluate: {orig_response} ### Feedback:
+Figure 3: Scoring prompt template adapted from [13], with red text indicating components that were perturbed in our experiments.
+We assess the scoring bias in LLM-as-a-Judge by perturbing the scoring prompt and measuring the resulting impact on judgments.
+Following [13], the input prompt $P$ for the scoring-based LLM judge consists of four components: a task description $T$, an evaluation target instance $I$, which can be either an instruction-response or a question-answer pair, reference answer(s) with corresponding score(s) $A$, and score rubrics $R$ as shown in Figure 3.
+The input prompt $P$ is formulated as follows:
+$P=(T,I^{*},[A],R),$ (1)
+where * denotes the component will not be perturbed, and $\cdot$] means that the enclosed component is optional within the input prompt.
+This formulation is central to our study.
+To specifically explore biases induced by non-target components (i.e., $T$, $R$, and $[A]$), we hold the evaluation target instance $I$ constant, as denoted by $I^{*}$.
+We also treat the reference answer $[A]$ as an optional element to explore its impact on scoring robustness, as it is not a mandatory part of the scoring process.
+By perturbing the reference answer $A\rightarrow A^{\prime}$ or the score rubrics $R\rightarrow R^{\prime}$, we get a perturbed input prompt $P^{\prime}$.
+At the same time, the task description $T$ will also be modified to $T^{\prime}$ in conjunction with the perturbed components to ensure the prompt remains coherent.
+The score from the perturbed prompt will be:
+$s^{\prime}=\textbf{LLM}(P^{\prime}).$ (2)
+For each target instance $I$, we have a golden score $s$ from advanced LLMs or human annotators.
+If $s$ and $s^{\prime}$ differ, it demonstrates that the LLM judge exhibits scoring bias on the given sample.
+3.2 Perturbation Types
+We propose three methods to perturb key parts of the judge prompt to introduce scoring biases:
+  • Score Rubric Order: The ordering of the description for each score within the rubric.
+We explore Ascending-Numeric (score 1 to score 5), Descending-Numeric (score 5 to score 1), and Random-Numeric orders (random score rubrics listed in prompt).
+  • Score ID: Typically, scores are represented using Arabic numerals $\{1,2,3,4,5\}$ as IDs.
+In this work, we explore the use of Letter-Grades $\{\text{E},\text{D},\text{C},\text{B},\text{A}\}$ and Roman-Numerals $\{\text{i},\text{ii},\text{iii},\text{iv},\text{v}\}$ as scoring IDs.
+  • Reference Answer Score: Based on the finding from [11, 13] that reference answers improve scoring accuracy for LLM judges, our work addresses an unexplored question: whether attaching a specific score $k$ to these references, denoted as Ref-$k$, could degrade scoring robustness.
+3.3 Datasets
+Following previous work [4, 13], we select four LLM-as-a-Judge datasets for scoring bias evaluation:
+  • BiGGen Bench [12]: This is a principled generation benchmark with instance-specific evaluation criteria.
+We use the non-multilingual portion of its test set, which contains 2,780 responses for 695 instances, all evaluated by human annotators.
+  • FLASK [48]: This is a fine-grained evaluation benchmark with 200 test prompts, 12 score rubrics, and 2,001 responses. It also includes scores labeled by human annotators.
+  • MT Bench [51]: This is a multi-turn chat benchmark with 80 test prompts.
+[11] write 80 hand-crafted score rubrics and generate 320 responses from WizardLM-13B [44], Vicuna-13B [5], Llama-2-Chat-13B [38], and GPT-3.5-Turbo-0613 [25].
+  • Vicuna Bench [5]: This is a single-turn chat benchmark with 80 test prompts.[11] also write 80 hand-crafted score rubrics and generate 320 responses from WizardLM-13B, Vicuna-13B, Llama-2-Chat-13B, and GPT-3.5-Turbo-0613.
+Table 1: Statistics of our evaluation datasets.
 Dataset Data Size Human Judgment Reference Answer
@@
 Vicuna Bench 320
-Table 1: Statistics of our evaluation datasets.
-Problem Formulation
-We assess the scoring bias in LLM-as-a-Judge by perturbing the scoring prompt and measuring the resulting impact on judgments.
-Following (Kim et al.
-2024c), the input prompt $P$ for the scoring-based LLM judge consists of four components: a task description $T$, a target instance $C^{*}$, which can be either an instruction-response or a question-answer pair, a reference answer with a fixed score $R$, and score rubrics $S$, as shown in Figure 3.
-The input prompt $P$ is formulated as follows:
-$P=(T,C^{*},[R],S),$ (1)
-where * denotes the component will not be perturbed, and $\cdot$] means that the enclosed component is optional within the input prompt.
-This work focuses on biases induced by perturbations on the non-target components of the input prompt for LLM judges.
-Accordingly, we keep the target instance fixed.
-We also treat the reference answer as an optional element to explore its impact on scoring robustness, as it is not a mandatory part of the scoring process.
-By perturbing the reference answer $R\rightarrow R^{\prime}$ or the score rubrics $S\rightarrow S^{\prime}$, we get a perturbed input prompt $P^{\prime}$.
-At the same time, the task description $T$ will also be modified to $T^{\prime}$ in conjunction with the perturbed components to ensure the prompt remains coherent.
-The score from the perturbed prompt will be:
-$s^{\prime}=\textbf{LLM}(P^{\prime}).$ (2)
-For each target instance $C$, we have a golden score $s$ from advanced LLMs or human annotators.
-If $s$ and $s^{\prime}$ differ, it demonstrates that the LLM judge exhibits scoring bias on the given sample.
-Perturbation Types
-We propose three methods to perturb key parts of the judge prompt to introduce scoring biases:
-  • Score Rubric Order: The ordering of the description for each score within the score rubrics. e.g., ascending order (score 1 to score 5), descending order (score 5 to score 1), or random order.
-  • Score ID: Typically, scores are represented using Arabic numerals $\{1,2,3,4,5\}$ as IDs.
-In this work, we explore the use of letter grades $\{\text{E},\text{D},\text{C},\text{B},\text{A}\}$ and Roman numerals $\{\text{i},\text{ii},\text{iii},\text{iv},\text{v}\}$ as scoring IDs.
-  • Reference Answer Score: Based on the finding from (Kim et al.
-2024a, c) that reference answers improve scoring accuracy for LLM judges, our work addresses an unexplored question: whether attaching specific scores to these references could degrade scoring robustness.
-Datasets
-Following previous work (Kim et al. 2024c; Chiang, Lee, and Lukasik 2025), we select four LLM-as-a-Judge datasets for scoring bias evaluation:
-  • BiGGen Bench (Kim et al.
-2024b): This is a principled generation benchmark with instance-specific evaluation criteria.
-We use the non-multilingual portion of its test set, which contains 2,780 responses for 695 instances, all evaluated by human annotators.
-  • FLASK (Ye et al. 2024b): This is a fine-grained evaluation benchmark with 200 test prompts, 12 score rubrics, and 2,001 responses. It also includes scores labeled by human annotators.
-  • MT Bench (Zheng et al.
-2023): This is a multi-turn chat benchmark with 80 test prompts.
-Kim et al.
-write 80 hand-crafted score rubrics and generate 320 responses from WizardLM-13B (Xu et al.
-2024), Vicuna-13B (Chiang et al.
-2023), Llama-2-Chat-13B (Touvron et al.
-2023), and GPT-3.5-Turbo-0613 (OpenAI 2023a).
-  • Vicuna Bench (Chiang et al.
-2023): This is a single-turn chat benchmark with 80 test prompts.Kim et al.
-also write 80 hand-crafted score rubrics and generate 320 responses from WizardLM-13B, Vicuna-13B, Llama-2-Chat-13B, and GPT-3.5-Turbo-0613.
-Model BiGGen Bench FLASK
+Table 2: Comparison of the alignment between our LLM judgment pipeline and human golden scores, where $\rho$ and $r$ represent Spearman’s and Pearson’s correlation coefficients, respectively.
+Model (Source of score) BiGGen Bench FLASK
 $\rho$ $r$ $\rho$ $r$
-GPT-4.1 0.6048 0.6360 0.6401 0.6823
-GPT-4o 0.6047 0.6340 0.6340 0.6557
-GPT-4 0.5741 0.6045 0.6207 0.6862
-Table 2: Comparison of the alignment between our LLM judgment pipeline and human golden scores, measured by Spearman’s $\rho$ and Pearson’s $r$ correlation coefficients.
-Scores for GPT-4.1 and GPT-4o are generated using our pipeline, whereas GPT-4 scores are sourced from the original benchmark.
+GPT-4.1 (our pipeline) 0.6048 0.6360 0.6401 0.6823
+GPT-4o (our pipeline) 0.6047 0.6340 0.6340 0.6557
+GPT-4 (original benchmark) 0.5741 0.6045 0.6207 0.6862
 The detailed statistics of our selected datasets are shown in Table 1.
-Due to the cost of manual annotation, the scores from advanced LLMs are usually used as golden scores (Ye et al.
-2024b; Kim et al.
-2024b).
-The instances in the above four datasets are pre-annotated with scores from GPT-4 (OpenAI 2023b).
-Due to the earlier release of GPT-4, we use more recent models GPT-4.1 (OpenAI 2025) and GPT-4o (OpenAI 2024) to obtain golden scores.
-Inspired by (Kim et al.
-2024a, c), we repeat the scoring process three times using a judge prompt that includes a reference answer assigned a score of 5, and then take a majority vote to obtain the final score from the LLM evaluators.
+Due to the cost of manual annotation, the scores from advanced LLMs are usually used as golden scores [12, 48].
+The instances in the above four datasets are pre-annotated with scores from GPT-4 [26], which is out of date at the current time.
+Therefore, we use more recent models such as GPT-4.1 [28] and GPT-4o [27] to obtain golden scores.
+Inspired by [11, 13], we repeat the scoring process three times using a judge prompt that includes a reference answer assigned a score of 5, and then take a majority vote to obtain the final score from the LLM evaluators.
 Table 2 shows that the GPT-4.1 scores exhibit the highest correlation with human scores.
@@
 To measure the impact of the reference answer score on the judgment robustness, it is necessary to additionally generate reference answers with scores ranging from 1 to 4.
-We propose a generation-review pipeline to synthesize reference answers for each sample in FLASK, as shown in Figure 4.
+We propose a generation-review pipeline to synthesize reference answers for each sample in BiGGen Bench and FLASK, as shown in Figure 4.
 Given an instruction and score rubrics, the response model generates the response according to the description from score 1 to 4.
@@
 In this pipeline, we use GPT-4.1 and GPT-4o, alternating their roles as the response and review models.
-Metrics
+3.4 Metrics
+For a dataset of $n$ samples, the human-annotated golden label set is denoted as $S=\{s_{1},\cdots,s_{n}\}$.
+We consider the ascending-numeric rubric without a reference answer as the baseline condition.
+The LLM outputs under this condition is denoted as $S^{(0)}=\{s_{1}^{(0)},\cdots,s_{n}^{(0)}\}$.
+Let $\mathcal{P}$ be the set of all perturbed prompt templates and the unperturbed baseline, and $\mathcal{P}^{\prime}$ be the set of all perturbed prompt templates excluding the baseline.
+For any prompt template $p$ in $\mathcal{P}$ or $\mathcal{P}^{\prime}$, the resulting LLM score set is $S^{(p)}=\{s_{1}^{(p)},\cdots,s_{n}^{(p)}\}$.
+Stability metrics quantify the consistency between the scores generated by a perturbed prompt $p,p\in\mathcal{P}^{\prime}$ (i.e.
+$S^{(p)}$) and the baseline scores $S^{(0)}$.
+We adopt flip rate (FP) and mean absolute deviation (MAD) as metrics.
+FP and MAD are defined as follows,
+$\text{FP}^{(p)}=\frac{1}{n}\sum_{i=1}^{n}\mathbb{I}(s_{i}^{(p)}\neq s_{i}^{(0)}),\quad\text{MAD}^{(p)}=\frac{1}{n}\sum_{i=1}^{n}|s_{i}^{(p)}-s_{i}^{(0)}|,\quad p\in\mathcal{P}^{\prime}.$ (3)
+Accuracy metrics quantify the agreement between the scores generated by a perturbed prompt $p,p\in\mathcal{P}$ (i.e.
+$S^{(p)}$) and the golden scores $S$.
+Specifically, we report both Spearman’s correlation coefficient $\rho$ [35] and Pearson’s correlation coefficient $r$ [30].
+Scoring tendency quantifies the model’s tendency to favor specific score categories for a given prompt template $p$ ($p\in\mathcal{P}$).
+For rating scale $\{r_{1},\cdots,r_{5}\}$ (e.g., $\{1,\cdots,5\}$), the tendency is defined by the following scoring distribution,
+$N^{(p)}=(N_{r_{1}}^{(p)},\cdots,N_{r_{5}}^{(p)}),\quad N_{r_{k}}^{(p)}=\sum_{i=1}^{n}{\mathbb{I}(s_{i}^{(p)}=r_{k})},\quad k\in\{1,\cdots,5\}.$ (4)
 Figure 4: The proposed generation-review pipeline for synthesizing reference answers.
-We quantify the effects of scoring bias on both scoring accuracy and scoring tendency.
-Given the golden scores, $S=\{s_{1},\cdots,s_{n}\}$, and the LLM judge’s scores, $S^{\prime}=\{s^{\prime}_{1},\cdots,s^{\prime}_{n}\}$, we quantify scoring accuracy as the correlation between $S$ and $S^{\prime}$.
-Specifically, we report both Spearman’s correlation coefficient $\rho$ (Spearman 1904) and Pearson’s correlation coefficient $r$ (Pearson 1895).
-The scoring tendency is quantified by the score distribution:
-$N_{k}$ $=\sum_{i=1}^{n}{\mathbb{I}(s_{i}=k)},\quad k\in\{1,2,3,4,5\},$ (3)
-$N$ $=(N_{1},N_{2},N_{3},N_{4},N_{5}).$ (4)
-Here, $N_{k}$ represents the count of score $k$ in the scoring results, and $N$ denotes the distribution of the score counts.
-Experiments
-Model Perturbation Strategy BiGGen Bench FLASK MT Bench Vicuna Bench
-Human GPT-4.1 Human GPT-4.1 GPT-4.1 GPT-4.1
-$\rho$ $r$ $\rho$ $r$ $\rho$ $r$ $\rho$ $r$ $\rho$ $r$ $\rho$ $r$
-GPT-4o Baseline - 0.5685 0.6165 0.7572 0.8059 0.5529 0.5914 0.7369 0.7676 0.7578 0.7724 0.6862 0.7696
-Rubrics Order Descending 0.5852 0.6245 0.7787 0.8233 0.5754 0.6080 0.7602 0.7760 0.7479 0.7597 0.6734 0.7635
-Random 0.5963 0.6313 0.7698 0.8124 0.5592 0.5893 0.7420 0.7573 0.7673 0.7838 0.6701 0.7504
-Score IDs Letter 0.5641 0.6034 0.7595 0.7994 0.5558 0.5977 0.7341 0.7609 0.7317 0.7454 0.6539 0.7712
-Roman numerals 0.5754 0.6171 0.7709 0.8148 0.5535 0.5912 0.7380 0.7641 0.7725 0.7786 0.7073 0.7821
-DeepSeek-V3 Baseline - 0.5314 0.5748 0.6775 0.7394 0.5163 0.5660 0.6711 0.7063 0.6323 0.6593 0.4313 0.5443
-Rubrics Order Descending 0.5356 0.5812 0.6611 0.7211 0.5181 0.5554 0.6555 0.6858 0.6183 0.6436 0.4993 0.5854
-Random 0.5136 0.5609 0.6547 0.7125 0.4999 0.5470 0.6521 0.6752 0.6178 0.6460 0.5086 0.5482
-Score IDs Letter 0.5378 0.5689 0.6803 0.7179 0.5314 0.5755 0.6663 0.6999 0.6489 0.6715 0.4832 0.5657
-Roman numerals 0.5164 0.5636 0.6688 0.7195 0.5066 0.5542 0.6750 0.7109 0.6558 0.6642 0.4416 0.5285
-Qwen3-32B Baseline - 0.5421 0.5855 0.7328 0.7715 0.5323 0.5695 0.6549 0.6807 0.6959 0.7053 0.6419 0.7188
-Rubrics Order Descending 0.5572 0.5989 0.7443 0.7862 0.5345 0.5657 0.6586 0.6839 0.6843 0.6989 0.6205 0.7020
-Random 0.5528 0.5938 0.7404 0.7824 0.5115 0.5456 0.6537 0.6748 0.6554 0.6688 0.6416 0.7252
-Score IDs Letter 0.5353 0.5851 0.7280 0.7770 0.5393 0.5724 0.6650 0.6975 0.6795 0.6930 0.6157 0.6808
-Roman numerals 0.5139 0.5629 0.6992 0.7412 0.5257 0.5781 0.6630 0.6980 0.6360 0.6418 0.6072 0.7126
-Qwen3-8B Baseline - 0.5092 0.5308 0.6522 0.6492 0.5035 0.5309 0.6464 0.6616 0.6238 0.6342 0.6639 0.7284
-Rubrics Order Descending 0.5241 0.5456 0.6458 0.6559 0.4885 0.5339 0.6315 0.6566 0.6005 0.6135 0.5500 0.6647
-Random 0.4869 0.5025 0.6091 0.5950 0.4778 0.5004 0.6061 0.6216 0.5949 0.5983 0.5135 0.6053
-Score IDs Letter 0.4853 0.5117 0.6410 0.6665 0.4873 0.5244 0.6334 0.6586 0.5985 0.6126 0.5812 0.6890
-Roman numerals 0.4801 0.5120 0.6231 0.6445 0.4606 0.5087 0.6177 0.6468 0.6016 0.5975 0.6111 0.7178
-Mistral-24B Baseline - 0.5059 0.5495 0.6910 0.7201 0.4950 0.5385 0.6303 0.6593 0.5943 0.5762 0.6116 0.6917
-Rubrics Order Descending 0.5369 0.5761 0.6987 0.7213 0.4922 0.5327 0.6362 0.6534 0.6104 0.6102 0.6137 0.7082
-Random 0.5053 0.5450 0.6569 0.6841 0.4756 0.5042 0.6212 0.6338 0.5603 0.5602 0.5461 0.6418
-Score IDs Letter 0.5114 0.5632 0.6685 0.7084 0.4880 0.5339 0.6311 0.6594 0.5375 0.5485 0.5558 0.6608
-Roman numerals 0.5102 0.5590 0.6640 0.6962 0.4957 0.5494 0.6533 0.6872 0.5863 0.5907 0.5903 0.6905
-Table 3: Correlation between golden scores and judge model scores under various prompt perturbations.
-For each judge model, the first row (Baseline) presents the metrics for the original, unperturbed prompt.
-The subsequent rows show the results under different types of perturbations applied to the rubrics order and score IDs.
-The degradation in correlation scores compared to the baseline reflects the model’s scoring bias, indicating its lack of robustness to these perturbations.
-Model names are abbreviated: DeepSeek-V3 denotes DeepSeek-V3-67B and Mistral-24B denotes Mistral-Small-24B-Instruct.
-We report Spearman’s ($\rho$) and Pearson’s ($r$) correlation coefficient.
-Setup
+4 Experiments
+4.1 Setup
 Judge Models.
-Previous work (Gao et al.
-2024b; Li et al.
-2024c; Liu et al.
-2024b) has reached a consensus that judge models should be stronger models to ensure evaluation reliability (Ye et al.
-2024a).
+Previous work [7, 18, 21] has reached a consensus that judge models should be stronger models to ensure evaluation reliability [47].
 After preliminary experiments, we found that scoring bias is a common phenomenon in judge models based on various LLMs.
@@
 Therefore, we select diverse judge models that include closed-source, open-source, large-size, and small-size models.
-They are GPT-4o, DeepSeek-V3-671B (Liu et al.
-2024a), Qwen3-32B, Qwen3-8B (Yang et al.
-2025), and Mistral-Small-24B-Instruct-2501 (MistralAI 2025).
-These judge models we selected differ from the models used to generate responses in the evaluation datasets, thus mitigating potential self-enhancement bias (Zheng et al.
-2023; Ye et al.
-2024a).
+They are GPT-4o, DeepSeek-V3-671B [20], Qwen3-32B, Qwen3-8B [46], and Mistral-Small-24B-Instruct-2501 [24].
+These judge models we selected differ from the models used to generate responses in the evaluation datasets, thus mitigating potential self-enhancement bias [47, 51].
 Judge Prompt.
-The prompt we use for scoring-based judgment, adapted from (Kim et al.
-2024c), is shown in Figure 3.
+The prompt we use for scoring-based judgment, adapted from [13], is shown in Figure 3.
 The unperturbed prompt is the one with ascending score rubrics, Arabic numeral score IDs, and no reference answer.
 Hyper-parameters. To ensure a deterministic scoring process, which is crucial for a fair evaluation of bias, we set the temperature of the LLM judge to 0 during all scoring tasks.
-Main Results
-Figure 5: Spearman’s correlation between human annotator scores and judge model scores on the FLASK dataset, under perturbations applied to the reference answer scores.
+4.2 Main Results
+Table 3: Stability metrics FP (%) and MAD, comparing baseline scores with scores under various prompt perturbations.
+A higher metric indicates a greater difference from the baseline score.
+The perturbation method Ref-5 was only applied to BiGGen Bench and FLASK, as the other two datasets do not provide reference answers.
+Model Perturbation BiGGen Bench FLASK MT Bench Vicuna Bench
+FR MAD FR MAD FR MAD FR MAD
+GPT-4o Descending-Numeric 23.63 0.2885 23.44 0.2769 23.13 0.2594 20.00 0.2094
+Random-Numeric 22.55 0.2651 24.94 0.3023 28.44 0.3375 19.38 0.1938
+Letter-Grades 18.17 0.2122 18.24 0.2259 19.69 0.2188 19.38 0.2031
+Roman-Numerals 16.51 0.1892 15.94 0.2014 23.13 0.2688 15.63 0.1625
+Ref-5 45.54 0.5604 45.58 0.7166 - - - -
+DeepSeek-V3-671B Descending-Numeric 24.21 0.2867 21.89 0.2734 25.00 0.2969 27.50 0.2906
+Random-Numeric 30.86 0.3565 32.83 0.4123 32.19 0.3875 39.06 0.4250
+Letter-Grades 19.86 0.2349 18.14 0.2374 17.81 0.1969 26.25 0.2844
+Roman-Numerals 19.53 0.2338 17.89 0.2169 20.00 0.2344 25.00 0.2625
+Ref-5 38.99 0.4863 45.13 0.6742 - - - -
+Qwen3-32B Descending-Numeric 28.56 0.3281 25.54 0.2989 29.38 0.3656 34.06 0.3688
+Random-Numeric 26.12 0.2957 29.64 0.3448 33.44 0.3938 27.81 0.2906
+Letter-Grades 25.11 0.2770 24.34 0.2909 28.44 0.3313 29.06 0.3000
+Roman-Numerals 28.92 0.3259 28.39 0.3273 51.88 0.6875 30.94 0.3156
+Ref-5 37.59 0.7674 44.58 0.6512 - - - -
+Qwen3-8B Descending-Numeric 46.22 0.5296 32.23 0.3778 40.31 0.4594 44.06 0.4750
+Random-Numeric 34.93 0.3989 36.83 0.4363 35.31 0.4125 34.38 0.3750
+Letter-Grades 30.54 0.3493 28.39 0.3228 36.88 0.4063 28.44 0.2938
+Roman-Numerals 32.99 0.3827 34.38 0.4058 40.31 0.4344 30.94 0.3156
+Ref-5 36.48 0.4367 40.53 0.5712 - - - -
+Mistral-Small-24B-Instruct Descending-Numeric 31.04 0.3421 27.39 0.2904 32.81 0.3406 25.63 0.2719
+Random-Numeric 28.02 0.3083 33.03 0.3603 29.38 0.3066 22.81 0.2406
+Letter-Grades 25.36 0.2734 25.54 0.2809 26.88 0.2969 23.75 0.2438
+Roman-Numerals 21.83 0.2349 23.24 0.2549 21.56 0.2281 16.56 0.1688
+Ref-5 35.40 0.4065 48.58 0.6247 - - - -
+Table 4: Accuracy metrics Spearman’s ($\rho$) and Pearson’s ($r$) correlation coefficient, showing the correlation between judge scores under perturbations and golden scores.
+Each judge model’s first row serves as the unperturbed baseline.
+Results are colored green (better than baseline) or red (worse).
+Model Perturbation Human Score GPT-4.1 Score
+BiGGen Bench FLASK MT Bench Vicuna Bench
+$\rho$ $r$ $\rho$ $r$ $\rho$ $r$ $\rho$ $r$
+GPT-4o - 0.5669 0.6147 0.5529 0.5915 0.7578 0.7724 0.6862 0.7696
+[…]
+[…]
+[…]
+[…]
+DeepSeek-V3-671B - 0.5340 0.5823 0.5193 0.5746 0.6323 0.6593 0.4411 0.6119
+[…]
+[…]
+[…]
+[…]
+Qwen3-32B - 0.5418 0.5848 0.5323 0.5695 0.6959 0.7053 0.6419 0.7188
+[…]
+[…]
+[…]
+[…]
+Qwen3-8B - 0.5087 0.5419 0.5035 0.5309 0.6238 0.6342 0.6639 0.7284
+[…]
+[…]
+[…]
+[…]
+Mistral-Small-24B-Instruct - 0.5067 0.5543 0.4948 0.5392 0.5943 0.5762 0.6116 0.6917
+[…]
+[…]
+[…]
+[…]
 The existence of scoring bias is proved in all judge models.
-Table 3 and Figure 5 show the correlation between judge model scores and golden scores under all perturbation methods.
+Table 3 shows the difference between baseline model scores and perturbed model scores, Table 4 and Figure 6 show the correlation between judge model scores and golden scores under all perturbation methods.
 For all judge models we evaluated, including advanced models such as GPT-4o, we observed fluctuations in the scoring results under different types of perturbations.
@@
 (1) More powerful judge models tend to be more robust to scoring bias.
-Consistent with existing consensus (Gao et al.
-2024b; Li et al.
-2024c; Liu et al.
-2024b), our experimental results indicate that judge models with advanced capability and larger size tend to exhibit higher reliability in scoring.
-As shown in Table 3, for example, across all perturbation methods and datasets, the fluctuation in correlation coefficients between GPT-4o and the golden scores remains within 0.03, except for the case involving a reference answer with a score of 5.
-A comparison of the results for Qwen3-32B and Qwen3-8B shows that judge models with larger parameter sizes demonstrate greater robustness in scoring.
-In comparison with the GPT-4.1 scores in BiGGen-Bench and Vicuna Bench, the correlation coefficients of the Qwen3-8B scores exhibit maximum fluctuations exceeding 0.05 under score rubric order bias and score ID bias.
-With reference answer score bias, the maximum fluctuation reaches nearly 0.2.
-However, the changes in the correlation coefficient of Qwen3-32B under biases are generally within 0.03.
+Consistent with existing consensus [7, 18, 21], our experimental results indicate that judge models with advanced capability and larger size tend to exhibit higher reliability in scoring.
+As shown in Table 3, GPT-4o has FRs less than 25% and MADs less than 0.3 under perturbation in most cases, except for the case involving a full-marked reference answer.
+Althogh Qwen3-8B, has higher FRs and MADs than GPT-4o and the larger version Qwen3-32B, especially reaches 46.22% FR and 0.5296 MAD on BiGGen Bench under descending-numeric perturbation, which means that nearly half of the scores are corrupted.
 Therefore, more capable judge models should be prioritized in scoring evaluations.
@@
 (3) Judge models have their own scoring tendencies.
-Figure 6 presents the scoring tendencies of different judge models under various scoring biases on the FLASK dataset in the form of heatmaps.
+Figure 5 presents the scoring tendencies of different judge models under various scoring biases in the form of heatmaps.
 It is obvious that each judge model exhibits its own specific scoring preferences.
@@
 Other judge models show smoother scoring distribution shifts, but certain scores are still preferred.
-GPT-4o demonstrates a preference for a score of 4, while Qwen3-32B favors a score of 5.
+For example, GPT-4o demonstrates a preference for a score of 4.
 An examination of Figure 5 reveals that score rubric order bias and score ID bias have a certain impact on the scoring tendencies of the judge models; however, they do not fundamentally alter the overall score distribution.
@@
 Therefore, debiasing the judge models is crucial to ensuring the fairness and reliability of the evaluation.
-Figure 6: Score distributions of LLM judges under various prompt perturbations on the FLASK dataset.
-The horizontal axis represents the assigned score (where -1 indicates a failed judgment), and the vertical axis lists the different perturbation methods applied.
-For reference, the top row in each heatmap (Human) shows the distribution of human scores.
-The second row (None) shows the model’s score distribution for the original, unperturbed prompt.
-Methods for Scoring Bias Mitigation
-For advanced or large size judge models, applying scoring prompt that deviate from conventional preferences is worth considering.
-In almost all scoring-based evaluation tasks, Arabic numerals are commonly used as scoring IDs.
-The score rubric order typically follows a sequential order from score 1 to score 5.
-These choices align with human writing conventions and cognitive preferences in the design of scoring prompts.
-However, our experimental results shown in Table 3 indicate that the scoring prompts designed to align with the usual preferences do not necessarily produce the most accurate scoring results.
-As mentioned earlier, the performance of the GPT-4o judge model in scoring evaluations improves when Roman numerals are used as scoring IDs.
-In more than half of the cases, using letter grades as scoring IDs improves the scoring accuracy of DeepSeek-V3-671B.
-Regarding the score rubric order bias, adopting a descending order improves the scoring accuracy of GPT-4o, Qwen3-32B, and Mistral-Small-24B-Instruct in most instances.
+(a) BiGGen Bench
+(b) FLASK
+Figure 5: Score tendency of LLM judges on BiGGen Bench and FLASK, plotting score distributions (x-axis) against perturbation methods (y-axis).
+The “Human” (top) and “None” (second) rows serve as baselines, showing human scores and the model’s unperturbed scores, respectively.
+4.3 Methods for Scoring Bias Mitigation
+4.3.1 (1) Leveraging Model Scale as a Debiasing Strategy.
+Our findings show a strong correlation between model scale and bias resilience.
+As shown in Table 3, state-of-the-art models like GPT-4o consistently demonstrate high stability, with FRs often remaining below 25%.
+In stark contrast, smaller models are highly vulnerable.
+Qwen3-8B, for example, suffers a 46.22% FR on BiGGen Bench under a simple rubric reordering.
+The larger Qwen3-32B is far more stable (28.56% FR) under the same perturbation, confirming the link to model scale.
+This trend extends to accuracy, as shown in Table 4.
+Perturbations often improve GPT-4o’s correlations with human scores, whereas the same perturbations frequently degrade the accuracy of smaller models like Qwen3-8B and Mistral-Small.
+Therefore, for high-stakes evaluations, selecting a powerful model like GPT-4o is a primary strategy for ensuring robust and reliable results, justifying potential trade-offs in cost or latency.
+4.3.2 (2) Debiasing via Intentional Deviation from Convention.
+In most cases of scoring-based evaluation tasks, Arabic numerals are commonly used as scoring IDs, and the score rubric order typically follows a sequential order (1 to 5).
+These conventional choices align with human cognitive preferences in prompt design.
+However, our experimental results (Table 4) indicate that such prompts do not necessarily produce the most accurate scoring.
+For example, the GPT-4o judge improves when using Roman numerals.
+Besides, using letter grades as scoring IDs improves the scoring accuracy of DeepSeek-V3-671B.
+Regarding the score rubric order bias, adopting a descending order improves the scoring accuracy of DeepSeek-V3 and Mistral-Small-24B-Instruct in most instances.
 Intuitively, employing a random scoring rubric order tends to negatively impact the scoring accuracy of all judge models.
 Based on our analysis, we propose a core design principle for scoring prompts: intentional yet meaningful deviation from convention.
-Effective strategies include introducing structured, unconventional formats, such as presenting scoring rubrics in a descending order or using non-numeric identifiers like letter grades (A, B, C).
-However, the key is to maintain clarity and logical coherence.
-Therefore, chaotic perturbations—such as a fully randomized rubric order or non-hierarchical identifiers—should be avoided.
-Moreover, inspired by previous work (Liu et al.
-2024c; Park et al.
-2024), perturbations applied to scoring prompt also show potential as debiasing strategies.
-For example, effective approaches include implementing multiple variations of the scoring rubric order or IDs and then aggregating the results, or adversarially training the models on samples enhanced with scoring bias perturbations.
-Using full-mark reference answers can stabilize scoring preferences and achieve optimal accuracy.
-The impact of reference answer score bias on the scoring accuracy and scoring tendency of the judge model is presented in Figure 5 and Figure 6.
-The experimental results indicate that, compared to the score rubric order bias and the score ID bias, the influence of the reference answer score bias on scoring robustness is both more pronounced and more regular.
-Figure 5 illustrates the impact of introducing reference answers with different scores in the input prompt on the scoring accuracy of the judge model.
-For GPT-4o, DeepSeek-V3-671B, and Mistral-Small-24B-Instruct, the scoring accuracy in most cases improves when reference answers are introduced.
-However, for the two Qwen judge models, only reference answers with scores of 4 or 5 contribute to enhancing scoring performance.
-Figure 5 also reveals an interesting phenomenon: the higher the score of the reference answer, the better the scoring accuracy of the judge model.
-Although DeepSeek-V3-671B, Qwen3-32B, and Mistral-Small-24B-Instruct do not strictly adhere to this pattern, the correlation between their scores and human scores similarly exhibits an upward trend as the reference answer score increases.
-Figure 6 shows that the score of the reference answer has a significant impact on the scoring tendency of the judge model.
-Specifically, the scoring results from DeepSeek-V3-671B, Qwen3-32B, and Qwen3-8B appear to be highly influenced by the score of the reference answer.
-When the score of the reference answer is between 1 and 4, these models tend to assign scores similar to that of the reference answer.
-Interestingly, when a reference answer with a score of 5 is introduced, the scoring behavior of the judge model appears to become more rational.
-This is particularly notable for DeepSeek-V3-671B, which exhibits a strong tendency to assign a score of 5 without bias.
-However, the introduction of a reference answer scored at 5 mitigates this phenomenon.
-The above analysis reveals that the score of the reference answer introduced in the prompt has a significant impact on the judge model’s scoring behavior.
+Effective strategies include structured, unconventional formats, such as presenting rubrics in descending order or using non-numeric identifiers like letter grades (A, B, C).
+However, the key is to maintain clarity.
+Therefore, chaotic perturbations—such as a fully randomized rubric order—should be avoided as they harm logical coherence.
+Moreover, inspired by previous work [23, 29], perturbations applied to scoring prompt also show potential as debiasing strategies.
+4.3.3 (3) Using full-mark reference answers can stabilize scoring preferences and achieve optimal accuracy.
+The impact of reference answer score bias on scoring accuracy and tendency is shown in Figure 6 and Figure 5.
+Our results indicate this bias’s influence on robustness is more pronounced and regular than that of score rubric order or ID bias.
+As shown in Figure 6, introducing reference answers generally improves scoring accuracy for models like GPT-4o, DeepSeek-V3-671B, and Mistral-Small-24B-Instruct.
+Notably, across both BiGGen Bench and FLASK, introducing a full-marked (score 5) reference answer consistently yields top-tier scoring accuracy, significantly outperforming other scores and the baseline.
+Figure 5 reveals a significant impact on scoring tendency, especially for DeepSeek-V3-671B, Qwen3-32B, and Qwen3-8B.
+When the reference score is between 1 and 4, these models tend to assign similar scores, indicating a strong pull from the reference.
+Interestingly, when a score 5 reference answer is introduced, the judge’s scoring behavior becomes more rational.
 In particular, when reference answers with a score of 5 are introduced, the models achieve the highest scoring accuracy without disproportionately influencing their scoring preferences.
-Therefore, it can be concluded that when incorporating reference answers into a scoring-based judge prompt, a full-mark reference answer is the best choice.
-Interestingly, this conceptually aligns with the principles of traditional statistical metrics, such as BLEU and ROUGE, which emphasize the critical role of golden answers in evaluation.
-Limitations and Future Work
+Therefore, when incorporating reference answers into a scoring-based judge prompt, a full-mark reference answer is the best choice.
+Conceptually, this aligns with metrics like BLEU and ROUGE, which also rely on golden answers.
+(a) BiGGen Bench
+(b) FLASK
+Figure 6: Accuracy metric Spearman’s correlation between human scores and judge model scores on BiGGen Bench and FLASK dataset, under perturbations applied to the reference answer scores.
+5 Limitations and Future Work
 While our work offers notable conclusions and insights, several limitations remain to be addressed in future work.
@@
 Although there are no dedicated methods for addressing scoring bias, intuitive and straightforward approaches, such as scoring multiple times followed by majority voting or averaging the scores, require empirical validation for their effectiveness in mitigating scoring bias.
-Moreover, the elaborate design of scoring prompt templates, the construction of debiased scoring-based judgment datasets (Park et al.
-2024), and adversarial training are potential approaches to enhancing the scoring robustness of judge models.
+Moreover, the elaborate design of scoring prompt templates, the construction of debiased scoring-based judgment datasets [29], and adversarial training are potential approaches to enhancing the scoring robustness of judge models.
 Additionally, although our work demonstrates the existence of scoring bias in LLM-as-a-Judge, the underlying causes of scoring bias remain to be validated.
-Approaches such as training data analysis and information flow observation (Yuan et al.
-2025) may help identify the reasons for scoring bias, whether it originates from within the model or from external factors.
-Conclusion
+Approaches such as training data analysis and information flow observation [49] may help identify the reasons for scoring bias, whether it originates from within the model or from external factors.
+6 Conclusion
 This work investigates scoring bias in LLM-as-a-Judge, where scoring judgments shift due to minor perturbations in judge prompts.
-To systematically evaluate this, we designed a framework that perturbs score rubric order, score IDs, and reference answers, quantifying the bias using scoring accuracy and scoring tendency.
+To evaluate this, we designed a framework that perturbs score rubric order, score IDs, and reference answers, quantifying the bias using scoring stability, scoring accuracy and scoring tendency.
 Experiments on synthesized datasets expanded from four benchmarks confirm that scoring bias significantly undermines the robustness and reliability of existing LLM judges, with varying sensitivity across different models.
 To mitigate this, we recommend unconventional prompt designs—such as descending rubric order or using letter or Roman numeral IDs—and providing a full-mark reference answer when applicable, both of which effectively improve performance.
-In summary, our work provides systematic evidence and actionable recommendations for evaluating and enhancing the reliability of LLM judges.
+In summary, our work provides evidence and actionable recommendations for evaluating and enhancing the reliability of LLM judges.
+{credits}
+[…]
+The authors have no competing interests to declare that are relevant to the content of this article.